The difference between three times a number and 7 is 17
Answer by ankor@dixie-net.com(22740)

The difference between three times a number and 7 is 17
:
Just write an equation for what it says, and follow the algebra rules
3n - 7 = 17
3n = 17 + 7
3n = 24
n = 24/3
n = 8
:
:
Confirm it in the original equation
3(8) - 7 = 17
24 - 7 = 17